# Investigation of the Challenges of Underwater-Visual-Monocular-SLAM

### v1

@@ Abstract
In this paper, we present a comprehensive investigation of the challenges of Monocular Visual Simultaneous Localization and Mapping (vSLAM) methods for underwater robots. While significant progress has been made in state estimation methods that utilize visual data in the past decade, most evaluations have been limited to controlled indoor and urban environments, where impressive performance was demonstrated. However, these techniques have not been extensively tested in extremely challenging conditions, such as underwater scenarios where factors such as water and light conditions, robot path, and depth can greatly impact algorithm performance. Hence, our evaluation is conducted in real-world AUV scenarios as well as laboratory settings which provide precise external reference. A focus is laid on understanding the impact of environmental conditions, such as optical properties of the water and illumination scenarios, on the performance of monocular vSLAM methods. To this end, we first show that all methods perform very well in in-air settings and subsequently show the degradation of their performance in challenging underwater environments. The final goal of this study is to identify techniques that can improve accuracy and robustness of SLAM methods in such conditions. To achieve this goal, we investigate the potential of image enhancement techniques to improve the quality of input images used by the SLAM methods, specifically in low visibility and extreme lighting scenarios in scattering media. We present a first evaluation on calibration maneuvers and simple image restoration techniques to determine their ability to enable or enhance the performance of monocular SLAM methods in underwater environments.
@@ 1 Introduction
Figure 1: (Left) Sparse Colmap reconstruction of the real Girona 500 Series AUV A3 surveying mission. After an initialization maneuver, a typical lawn mower patter followed by two cross tracks to support loop closing attempts are executed. The camera trajectory is drawn in red. Left lower magnification: increased number of loop closing opportunities (correspondences drawn in Pink) by executing cross-tracks, while the upper one shows a top- and a side-view of the initialization maneuver carried out to support SLAM algorithms. (Right) AUV in deep underwater mission and deployed camera-light-system.
@@
In addition to monocular vSLAM methods, sensor fusion algorithms are crucial for accurate and robust navigation of underwater robots. Underwater environments pose unique challenges, such as low visibility and unpredictable terrain, that can significantly impact the accuracy of navigation systems. Therefore, integrating data from multiple sensors, such as Inertial Navigation Systems (INS), sonars, lasers, and cameras, can improve the overall performance of the navigation system. For example, combining data from an Inertial Navigation System (INS) and a Doppler Velocity Log (DVL) can improve the accuracy of underwater robot navigation by providing velocity measurements that are not affected by currents. Similarly, integrating data from an INS and a sonar can improve underwater localization by providing depth information that can be used to correct INS drift. Sonar sensors, such as multibeam sonars, can also be used to provide high-resolution 3D maps of underwater environments that can be used for localization and mapping [Drews Junior et al., 2016]. Laser-based sensors, such as scanning laser rangefinders, can also be used to provide high-resolution 3D maps of underwater environments that can be used for localization and mapping [Palomeras et al., 2016]. Incorporating data from multiple sensors can be challenging due to the different modalities and measurement noise associated with each sensor. However, recent advancements in sensor fusion algorithms, such as Extended Kalman Filters (EKF) and Unscented Kalman Filters (UKF), have made it possible to effectively combine data from multiple sensors in real-time [Yu et al., 2019] [Yang et al., 2019].
While sensor fusion using multiple modalities has shown great potential in improving underwater robot navigation and mapping, it also comes with additional cost and complexity. Therefore, in this paper, we evaluate the performance of offline and online monocular vSLAM methods for underwater robot navigation in both a real-world and a water tank in a laboratory setting. Our focus is on the impact of environmental conditions, such as water and luminance, on the performance of monocular vSLAM methods, to this end, we first show the good performance of the chosen SLAM methods in-air and then investigate their performance degradation with respect to varying environmental conditions in a scattering medium, i.e., water. Finally, we investigate the potential of image enhancement techniques to improve accuracy and robustness in challenging underwater environments.
@@ 2 Monocular vSLAM
Monocular vSLAM methods are commonly classified into three categories: feature-based methods, direct methods, and visual-inertial methods [Zhou et al., 2019]. Feature-based methods, such as ORB-SLAM2 [Mur-Artal et al., 2017], rely on detecting and tracking distinctive features in the image frames to estimate camera poses and create a map of the environment. These methods have been shown to achieve accurate results in a variety of scenarios, including indoor and outdoor environments. However, they can be sensitive to changes in illumination, texture, and occlusions, which can cause feature tracking failures and affect their robustness [Mur-Artal and Tardós, 2015]. A newer versions of ORB-SLAM2, ORB-SLAM3 [Campos et al., 2021] incorporate semantic information to improve the robustness and accuracy of the system. Direct methods, such as DSO [Engel et al., 2018] and LSD-SLAM [Engel et al., 2014] estimate camera motion and 3D structure directly from the intensity values of the image frames, without relying on feature detection and tracking. LSD-SLAM uses semi-dense depth maps to estimate the camera poses and create a map of the environment. This method is known for its ability to handle large-scale environments and low-texture scenes while ORB-SLAM2 uses ORB features to detect and track keypoints in the image frames, and then estimates the camera poses and creates a map of the environment based on the feature matches. BAD-SLAM [Bian et al., 2020] is another direct method that estimates the camera motion and 3D structure directly from the intensity values of the image frames. It aims to improve the accuracy and robustness of SLAM systems by directly leveraging the RGB-D information and performing real-time bundle adjustment.
These methods can provide accurate and dense reconstructions in low-texture environments, but they require significant computation power and are less robust to changes in illumination and scene geometry [Engel et al., 2017].Visual-inertial methods, such as OKVIS [Leutenegger et al., 2015] and VINS-Mono [Qin et al., 2018], fuse camera and inertial measurements to estimate camera poses and create a map of the environment. These methods can achieve accurate results in highly dynamic and challenging environments, but they require additional sensors and calibration [Forster et al., 2017]. However, these methods can face challenges when operating in highly dynamic underwater or low-light environments [Köser and Frese, 2020]. To address these challenges, researchers have proposed various modifications to existing methods or developed entirely new methods. For example, the authors of [Ferrera et al., 2019] proposed a new visual odometry method specifically designed to handle the challenging conditions of underwater environments without any previous image enhancement step. Howerver, in our study we focus only monocular vSLAM methods that use camera data alone and do not incorporate inertial measurements.
@@ 3 Image Enhancement / Restoration
Underwater environments typically exhibit extreme light conditions and poor visibility, which can significantly impair the performance of monocular vSLAM methods. To mitigate this issue, researchers have proposed various techniques for enhancing underwater images, which we here broadly divide into four categories: heuristic, statistical, physically-based, and machine learning methods (see [Song et al., 2022] for a comprehensive survey).
@@ 3.1 Statistical methods
Statistical methods for underwater image restoration aim to recover the original image from its degraded underwater version. These methods use statistical models to estimate the degradation factors, such as light attenuation, scattering, and noise, and then use these estimates to restore the image. For instance, [Chiang and Chen, 2012] proposed a wavelength compensation and dehazing approach that estimates the light attenuation coefficient and compensates for color distortion caused by the water medium. In [Ancuti et al., 2012] the authors proposed an underwater image enhancement method that relies only on the degraded version of the image for input and weight measures. They used two inputs to represent color correction and contrast enhancement of the original underwater image/frame, while four weight maps are employed to enhance the visibility of distant objects degraded by medium scattering and absorption. Drews Jr. et al. [Drews Jr et al., 2017] proposed a statistical approach that estimates the parameters of a degradation model and inverts the degradation process to restore the original image.
@@ 3.2 Heuristic methods
Heuristic-based methods for underwater image restoration exploit specific underwater environment characteristics. Li et al.’s underwater dark channel prior (UDCP) [Li et al., 2016] uses the dark channel prior principle to estimate transmission and restore the image. Kim and Lee’s adaptive histogram equalization (AHE) [Kim and Lee, 2017] applies histogram equalization to small image regions for contrast enhancement. Pizer et al.’s contrast-limited adaptive histogram equalization (CLAHE) [Pizer et al., 1987] limits contrast enhancement to prevent over-amplification of noise. In [Köser et al., 2021], the authors present a practical approach to compensating for these lighting effects on flat seafloor regions found in the Abyssal plains. The method is parameter-free and performs robust statistics-based estimates of additive and multiplicative nuisances without requiring explicit parameters for light, camera, water, or scene. Although heuristic models can produce impressive outcomes in situations that align with their underlying assumptions, their results may lack consistency in other scenarios. Moreover, they do not assure to adhere to physical principles.
@@ 3.3 Machine Learning methods
Machine learning-based methods have shown promising results in enhancing underwater images by learning from a large dataset of annotated images. One popular machine learning-based method is the Deep Underwater Image Enhancement (DUIE) framework proposed by Zhang et al. [Zhang et al., 2019]. DUIE uses a convolutional neural network (CNN) to learn the mapping between the low-quality input underwater image and the high-quality output image. Generative Adversarial Networks (GANs) are a type of deep neural network that consist of a generator and a discriminator. GANs have been used to learn the mapping between degraded and enhanced underwater images. The Underwater GAN (UW-GAN) proposed by Li et al. [Li et al., 2018] aims to improve the visibility of underwater images. UW-GAN uses an underwater image dataset to train the generator to generate enhanced versions of degraded underwater images. Other GAN-based methods for underwater image enhancement include the Conditional GAN (CGAN) [Fu et al., 2018] and the Multi-Scale GAN (MSGAN) [Xu et al., 2019]. Despite benefiting from the expressive capabilities of neural networks, machine learning methods are typically trained under specific conditions. However, underwater environments are often characterized by dynamic conditions and their unpredictability, which can pose challenges to these methods.
@@ 3.4 Physically-based methods
Physically-based methods for underwater image restoration aim to model the physically-based processes that cause image degradation, such as light attenuation, scattering, and absorption, and then invert these models to restore the image. In doing so, they are the only methods able to do image restoration as opposed to image enhancement. These methods typically require knowledge of the physical properties of parts of the scene and can be computationally intensive. [Garcia et al., 2017] proposed a graph-based algorithm for color correction of underwater images. This method uses a graph-based representation to model the relationships between different color channels in the image and to estimate the color correction factors. Another method is the Sea-thru method proposed by Akkaynak and Treibitz [Akkaynak and Treibitz, 2018] which estimates the backscatter using the dark pixels and their known range information, and then uses an estimate of the spatially varying illuminant to obtain the range-dependent attenuation coefficient. The latter method, however, is only valid for the Sun-light case, which exhibits homogeneous illumination. [Boittiaux et al., 2023] implemented multiview extensions, to improve the estimates and applied the method to datasets with artificial illumination. However, to make this approach work, the artificial illumination has to be locally homogeneous. Further approaches, which can be applied to true heterogeneous underwater artificial scenarios, i.e., with artificial illumination, are presented in [Bryson et al., 2016] and [Nakath et al., 2021]. While physically-based methods have the advantage of being based on well-established principles, they can be limited by the availability and accuracy of the precise parameters needed for the models.
@@ 4 Datasets
With Girona 500 series AUVs, we collected A-datasets in real waters, as well as T-datasets in a water tank with a precise ground truth estimate. In all settings, we carried out dedicated calibration maneuvers, which foster the initialization of SLAM approaches. Furthermore, in the tank and the AUV sets, we carried out classical lawn mower patterns with a stable flying height with subsequent cross-tracks, to support loop-closing approaches. In the tank, we additionally recorded sets, wich resemble a more free-flying scanning path. The latter brings about a lot of loop closing opportunities, which will however be impaired by big height variances, which in turn induce big changes in visual appearance in scattering media.
@@ 4.1 Water Test Tank with Ground Truth
We equipped a $2.2\times 1\times 0.8$m water test tank with three 50w Wasler daylight bulbs (5400k) housed in Walimex diffusors to create a homogeneous illumination setting akin to heavy atmospheric scattering. In addition, we attached two Ulanzi L2 Lite (5500k) co-moving lights, to be able to simulate active underwater light systems to a custom-build underwater camera.
After building a small-scale test scene, we took several sets, to acquire underwater imagery with external reference as ground truth. As this is close to impossible in real waters, we equipped the camera with a stick and attached two Vive controllers, whose pose (position and attitude) in space can be precisely determined in air. This information can be used, to obtain a fused estimate of the pose of the underwater camera. We found the mean accuracy of the system tracking performance to be smaller 3 mm and 0.3 deg for translation and attitude, respectively [Winkel et al., 2023]. For the dataset, an in-air fisheye calibration was conducted, with a residual error of 0.22px. Then, we center the camera within a dome port, to exclude refraction effects from the dataset, stemming from the traversal of interfaces of media with different optical densities [She et al., 2019, She et al., 2022]. Finally, we conducted an underwater fisheye calibration of the camera, with residual reprojection error of 0.55px to capture remaining disturbances, which have not been captured by the preceding steps. Hence, we will exclusively deal with color distortion effects in those datasets. Specifically, we took homogeneously illuminated sets (T1-3), sets with mixed illumination (T4/5) and finally two sets with co-moving lights (T6/7) in air (see Fig. 3). Subsequently, we added water and dye for the attenuation as well as Maaloxan as the scattering agent until the working range was clearly distorted by the corresponding effects. This setup enables us to mimic the underwater conditions in Sunlight (T8/9), a mixed (Sun-artificial) light scenario (T10/11), as well as deep sea conditions, where only the artificial light is visible (T12/13; see Fig. 3). All former sets of those pairs execute a lawn mover pattern, while all latter sets execute free scanning trajectories with bigger depth variances.

### v2

@@ Abstract
In this paper, we present a comprehensive investigation of the challenges of Monocular Visual Simultaneous Localization and Mapping (vSLAM) methods for underwater robots. While significant progress has been made in state estimation methods that utilize visual data in the past decade, most evaluations have been limited to controlled indoor and urban environments, where impressive performance was demonstrated. However, these techniques have not been extensively tested in extremely challenging conditions, such as underwater scenarios where factors such as water and light conditions, robot path, and depth can greatly impact algorithm performance. Hence, our evaluation is conducted in real-world AUV scenarios as well as laboratory settings which provide precise external reference. A focus is laid on understanding the impact of environmental conditions, such as optical properties of the water and illumination scenarios, on the performance of monocular vSLAM methods. To this end, we first show that all methods perform very well in air and subsequently investigate the degradation of their performance in ever more challenging underwater environments. The final goal of this study is to identify techniques that can improve accuracy and robustness of SLAM methods in such conditions. To achieve this goal, we investigate the potential of image enhancement techniques to improve the quality of input images used by the SLAM methods, specifically in low visibility and extreme lighting scenarios in scattering media. We present a first evaluation on calibration maneuvers and simple image restoration techniques to determine their ability to enable or enhance the performance of monocular SLAM methods in underwater environments.
@@ 1 Introduction
Figure 1: (Middle) Sparse Colmap reconstruction of the real Girona 500 Series AUV A3 surveying mission. After an initialization maneuver, a typical lawn mower patter followed by two cross tracks to support loop closing attempts are executed. The camera trajectory is drawn in red. Left lower magnification: increased number of loop closing opportunities (correspondences drawn in Pink) by executing cross-tracks, while the upper one shows a top- and a side-view of the initialization maneuver carried out to support SLAM algorithms. (Right) AUV in deep underwater mission and deployed camera-light-system.
@@
In addition to monocular vSLAM methods, sensor fusion algorithms are crucial for accurate and robust navigation of underwater robots. Underwater environments pose unique challenges, such as low visibility and unpredictable terrain, that can significantly impact the accuracy of navigation systems. Therefore, integrating data from multiple sensors, such as Inertial Navigation Systems (INS), sonars, lasers, and cameras, can improve the overall performance of the navigation system. For example, combining data from an Inertial Navigation System (INS) and a Doppler Velocity Log (DVL) can improve the accuracy of underwater robot navigation by providing velocity measurements that are not affected by currents. Similarly, integrating data from an INS and a sonar can improve underwater localization by providing depth information that can be used to correct INS drift. Sonar sensors, such as multibeam sonars, can also be used to provide high-resolution 3D maps of underwater environments that can be used for localization and mapping [Palomer et al., 2016]. Laser-based sensors, such as scanning laser rangefinders, can also be used to provide high-resolution 3D maps of underwater environments that can be used for localization and mapping [Vila, 2018]. Incorporating data from multiple sensors can be challenging due to the different modalities and measurement noise associated with each particular sensor. However, advancements in sensor fusion algorithms, such as Extended Kalman Filters (EKF) and Unscented Kalman Filters (UKF), have made it possible to effectively combine data from multiple sensors in real-time [Wu et al., 2019].
While sensor fusion using multiple modalities has shown great potential in improving underwater robot navigation and mapping, it also comes with additional cost and complexity. Therefore, in this paper, we evaluate the performance of offline and online monocular vSLAM methods for underwater robot navigation in both a real-world and a water tank in a laboratory setting. Our focus is on the impact of environmental conditions, such as water and illumination, on the performance of monocular vSLAM methods. To this end, we first show the good performance of the chosen SLAM methods in air and then investigate their performance degradation with respect to varying environmental conditions in a scattering medium, i.e., water. Finally, we investigate the potential of image enhancement techniques to improve accuracy and robustness in challenging underwater environments.
@@ 2 Monocular vSLAM
Monocular vSLAM methods are commonly classified into three categories: feature-based methods, direct methods, and visual-inertial methods [Zhang et al., 2022]. Feature-based methods, such as ORB-SLAM2 [Mur-Artal et al., 2017], rely on detecting and tracking distinctive features in the image frames to estimate camera poses and create a map of the environment. These methods have been shown to achieve accurate results in a variety of scenarios, including indoor and outdoor environments. However, they can be sensitive to changes in illumination, texture, and occlusions, which can cause feature tracking failures and affect their robustness [Mur-Artal and Tardós, 2015]. Newer versions of ORB-SLAM2, ORB-SLAM3 [Campos et al., 2021] incorporate semantic information to improve the robustness and accuracy of the system. Direct methods, such as DSO [Engel et al., 2018] and LSD-SLAM [Engel et al., 2014] estimate camera motion and 3D structure directly from the intensity values of the image frames, without relying on feature detection and tracking. LSD-SLAM uses semi-dense depth maps to estimate the camera poses and create a map of the environment. This method is known for its ability to handle large-scale environments and low-texture scenes while ORB-SLAM2 uses ORB features to detect and track keypoints in the image frames, and then estimates the camera poses and creates a map of the environment based on the feature matches. BAD-SLAM [Schops et al., 2019] is another direct method that estimates the camera motion and 3D structure directly from the intensity values of the image frames. It aims to improve the accuracy and robustness of SLAM systems by directly leveraging the RGB-D information and performing real-time bundle adjustment.
These methods can provide accurate and dense reconstructions in low-texture environments, but they require significant computational power and are less robust to changes in illumination and scene geometry [Engel et al., 2017].Visual-inertial methods, such as OKVIS [Leutenegger et al., 2015] and VINS-Mono [Qin et al., 2018], fuse camera and inertial measurements to estimate camera poses and create a map of the environment. These methods can achieve accurate results in highly dynamic and challenging environments, but they require additional sensors and calibration [Forster et al., 2017]. However, also these methods face challenges when operating in highly dynamic underwater or low-light environments [Köser and Frese, 2020]. To address these challenges, researchers have proposed various modifications to existing methods or developed entirely new methods. For example, [Ferrera et al., 2019] proposed a new visual odometry method specifically designed to handle the challenging conditions of underwater environments without any previous image enhancement step. However, in our study we focus only monocular vSLAM methods that use camera data alone and do not incorporate inertial measurements.
@@ 3 Image Enhancement / Restoration
Underwater environments typically exhibit extreme light conditions and poor visibility, which can significantly impair the performance of monocular vSLAM methods. To mitigate this issue, researchers have proposed various techniques for enhancing underwater images, which we here broadly divided into four categories: heuristic, statistical, physically-based, and machine learning methods (see [Song et al., 2022] for a comprehensive survey).
@@ 3.1 Statistical methods
Statistical methods for underwater image restoration aim to recover the original image from its degraded underwater version. These methods use statistical models to estimate the degradation factors, such as light attenuation, scattering, and noise, and then use these estimates to restore the image. For instance, [Chiang and Chen, 2012] proposed a wavelength compensation and dehazing approach that estimates the light attenuation coefficient and compensates for color distortion caused by the water medium. In [Ancuti et al., 2012] the authors proposed an underwater image enhancement method that relies only on the degraded version of the image for input and weight measures. They used two inputs to represent color correction and contrast enhancement of the original underwater image/frame, while four weight maps were employed to enhance the visibility of distant objects degraded by medium scattering and absorption. [Ponce-Hinestroza et al., 2016] proposed a statistical approach that estimates the parameters of a degradation model and inverts the degradation process to restore the original image.
@@ 3.2 Heuristic methods
Heuristic-based methods for underwater image restoration exploit specific underwater environment characteristics. The underwater dark channel prior (UDCP) [Drews et al., 2013] uses the dark channel prior principle to estimate transmission and restore the image. Mohan’s adaptive histogram equalization (AHE) [Mohan and Simon, 2020] applies histogram equalization to small image regions for contrast enhancement. Pizer et al.’s contrast-limited adaptive histogram equalization (CLAHE) [Pizer et al., 1987] limits contrast enhancement to prevent over-amplification of noise. In [Köser et al., 2021], the authors present a practical approach to compensating for these lighting effects on flat seafloor regions found in the Abyssal plains. The method is parameter-free and performs robust statistics-based estimates of additive and multiplicative nuisances without requiring explicit parameters for light, camera, water, or scene. Although heuristic models can produce impressive outcomes in situations that align with their underlying assumptions, their results may lack consistency in other scenarios. Moreover, they do not assure to adhere to physical principles.
@@ 3.3 Machine Learning methods
Machine learning-based methods have shown promising results in enhancing underwater images by learning from a large dataset of annotated images. One popular machine learning-based method is the Underwater-CNN (UWCNN) proposed in [Anwar et al., 2018]. UWCNN uses a convolutional neural network (CNN) to learn the mapping between the low-quality input underwater image and the high-quality output image. Generative Adversarial Networks (GANs) are a type of deep neural network that consists of a generator and a discriminator. GANs have been used to learn the mapping between degraded and enhanced underwater images. The UWGAN proposed by Li et al. [Li et al., 2018] aims to improve the visibility of underwater images. It uses an underwater image dataset to train the generator to generate enhanced versions of degraded underwater images. Other GAN-based methods for underwater image enhancement include the Conditional GAN (CGAN) [Yang et al., 2020] and the Multi-Scale GAN [Guo et al., 2020]. Despite benefiting from the expressive capabilities of neural networks, machine learning methods are typically trained under specific conditions. However, underwater environments are often characterized by dynamic conditions and their unpredictability, which can pose challenges to these methods.
@@ 3.4 Physically-based methods
Physically-based methods for underwater image restoration aim to model the physically-based processes that cause image degradation, such as light attenuation, scattering, and absorption, and then invert these models to restore the image. In doing so, they are the only methods able to do image restoration as opposed to image enhancement. These methods typically require knowledge of the physical properties of parts of the scene and can be computationally intensive. [Sharma and Varma, 2022] proposed a graph-based algorithm for color correction of underwater images. The technique models color relationships via a graph structure and applies root-filtering in graph spectral domains, enhancing high-frequency details for improved visual quality. Another method is the Sea-thru method proposed by Akkaynak and Treibitz [Akkaynak and Treibitz, 2019] which estimates the backscatter using the dark pixels and their known range information, and then uses an estimate of the spatially varying illuminant to obtain the range-dependent attenuation coefficient. The latter method, however, is only valid for the Sun-light case, which exhibits homogeneous illumination. [Boittiaux et al., 2023] implemented multiview extensions, to improve the estimates and applied the method to datasets with artificial illumination. However, to make this approach work, the artificial illumination has to be locally homogeneous. Further approaches, which can be applied to true heterogeneous underwater artificial scenarios, i.e., with artificial illumination, are presented in [Bryson et al., 2016] and [Nakath et al., 2021]. While physically-based methods have the advantage of being based on well-established principles, they can be limited by the availability and accuracy of the precise parameters needed for the models.
@@ 4 Datasets
With Girona 500 series AUVs, we collected A-datasets in real waters, as well as T-datasets in a water tank with a precise ground truth estimate. In all settings, we carried out dedicated calibration maneuvers, which foster the initialization of SLAM approaches. Furthermore, in the tank and the AUV sets, we carried out classical lawn mower patterns with a stable flying height with subsequent cross-tracks, to support loop-closing approaches. In the tank, we additionally recorded sets, which resemble a more free-flying scanning path. The latter brings about a lot of loop closing opportunities, which will however be impaired by big height variances, which in turn induce big changes in visual appearance in scattering media.
@@ 4.1 Water Test Tank with Ground Truth
We equipped a 2.2×1×0.82.2$\times$ 1$\times$ 0.82.2 × 1 × 0.8m water test tank with three 50w Wasler daylight bulbs (5400k) housed in Walimex diffusors to create a homogeneous illumination setting akin to heavy atmospheric scattering. In addition, we attached two Ulanzi L2 Lite (5500k) co-moving lights, to be able to simulate active underwater light systems to a custom-built externally-tracked underwater camera [Winkel et al., 2023].
After building a small-scale test scene, we took several sets, to acquire underwater imagery with external reference as ground truth. As this is close to impossible in real waters, we equipped the camera with a stick and attached two Vive controllers, whose pose (position and attitude) in space can be precisely determined in air. This information can be used, to obtain a fused estimate of the pose of the underwater camera. We found the mean accuracy of the system tracking performance to be smaller 3 mm and 0.3 deg for translation and attitude, respectively [Winkel et al., 2023]. For the dataset, an in-air fisheye calibration was conducted, with a residual error of 0.22px. Then, we center the camera within a dome port, to exclude refraction effects from the dataset, stemming from the traversal of interfaces of media with different optical densities [She et al., 2019, She et al., 2022]. Finally, we conducted an underwater fisheye calibration of the camera, with residual reprojection error of 0.55px to capture remaining disturbances, which have not been captured by the preceding steps. Hence, we will exclusively deal with color distortion effects in those datasets. Specifically, we took homogeneously illuminated sets (T1-3), sets with mixed illumination (T4/5) and finally two sets with co-moving lights (T6/7) in air (see Fig. 3). Subsequently, we added water and dye for the attenuation as well as Maaloxan as the scattering agent until the working range was clearly distorted by the corresponding effects. This setup enables us to mimic the underwater conditions in Sunlight (T8/9), a mixed (Sun-artificial) light scenario (T10/11), as well as deep sea conditions, where only the artificial light is visible (T12/13); see Fig. 3). All former sets of those pairs execute a lawn mover pattern, while all latter sets execute free scanning trajectories with bigger depth variances.